OpenStudy (anonymous):

Compute the period of a spring for Hooke's constant 50 N/m hanging from a hook above the surface of: a) the earth b) the moon

OpenStudy (experimentx):

The time period of harmonic motion due to Hook law is given by \[T=1/2\pi(k/m)^1/2\] for equal weight masses have to be different on moon and earth i.e. mass should be more on moon than on earth. if m1, g1 and m2, g2 be the masses and accn due to gravity on earth and moon respectively, then, \[m1*g1=m2*g2\] if T1 and T2 be the time period on Earth and Moon respectively, then \[T1/T2=(m2/m1)^1/2 = (g1/g2)^1/2\] since the acceleration due to gravity is more on Earth than on moon, you have the results. P.S. I am always "quite not sure".